# Generative AI models enable efficient and physically consistent sea-ice simulations

### v1

@@ B.2.2 Variational flow matching
An interesting case arises when we employ the model with a single step, mapping directly from $\tau=0$ to $\tau=1$: there, the latent state for $\tau=0$ contains no information about the prediction. Consequently, if a neural network is trained for that, it would learn to disregard the latent state as input and to target a mean prediction [23], similar to the deterministic model. In this case, the model effectively reduces to a deterministic prediction model.
B.2.3 Learnable uncertainty weighting
The weighting in the loss function for the deterministic model and flow matching is determined by the climatological standard deviation. However, each of the six predicted variables is subject to distinct physical processes, leading to varying degrees of predictability; for example, the velocities are generally easier to predict than the thickness, which might be not captured by their climatological weighting [75].
We account for this in a data-dependent weighting by altering the Gaussian assumption of variational flow matching, Equation (29), and making some covariances learnable. As a simplification and exploiting the mean-field approximation, we introduce a per-variable standard deviation that depends on the pseudo time, $\mathbf{s}(\tau)$, obtaining
$q_{\boldsymbol{\theta}}(\mathbf{z}_{1}\mid\mathbf{z}_{\tau})=\mathcal{N}\left(\widehat{\boldsymbol{\mu}}_{\boldsymbol{\theta}},\frac{1}{2}\Big{(}(1-\tau)\cdot\mathbf{s}(\tau)\Big{)}^{2}\mathbf{I}\right),$ (32)
@@ B.2.3 Learnable uncertainty weighting
$\mathcal{L}_{\textrm{VFM}}(\boldsymbol{\theta},\mathbf{s})=\mathbb{E}_{\tau,\mathbf{z}_{0},\mathcal{D}}\left[\frac{1}{2\cdot\big{(}\mathbf{s}(\tau)\big{)}^{2}}\lVert\mathbf{u}-\widehat{\mathbf{v}}_{\boldsymbol{\theta}}(\mathbf{z},\tau,\mathbf{x}_{t},\mathbf{F}_{t:t+12h})\rVert^{2}_{2}+\log\big{(}\mathbf{s}(\tau)\big{)}\right],$ (33)
which has a similar formulation for diffusion models [61]. The squared standard deviation, $\big{(}\mathbf{s}(\tau)\big{)}^{2}$, converges to half of the expected per-variable mean-squared error of the velocity and adapts automatically to the predictability of different variables.
Figure 9: Learned weighting of the flow matching loss (a) The learned weighting, averaged across all six variables, at different resolutions indicates the importance of difference scales at different resolutions. (b) Used in an importance-sampling-like mechanism, the learned standard deviation can guide the tuning of the pseudo time scheduler during prediction. Here, the focus is on smaller scales, as opposed to a uniform scheduler.
Note that, while the weighting in the loss function is learned, the output of the neural network is still scaled by the climatological standard deviation. This disentanglement between the scaling and the learned weighting becomes especially interesting when the neural network is trained across different datasets or resolutions: we can set a constant scaling of the neural network and learn different weightings per dataset (Fig. 9a). The weighting concentrates around the relevant scales, acting as gradient normalisation and stabilising the training [61].
This learnable weighting can reduce the need to tune the scheduler during training and prediction. We can also use the learned standard deviation to guide our inference scheduler (Fig. 9b). The larger the learned standard deviation, the larger the expected error for that scale, and importance sampling suggests that we should oversample these scales [76]. Hence, we want to concentrate our pseudo time scheduler during prediction to the scales where the expected error is large, which are in our case the small scales. While these small scales might be unimportant for perceptual generation like images, they can become especially important when the model is auto-regressively cycled, as errors at small scales can amplify over time and impact larger scales. As prediction scheduler, we use a shifted sigmoid weighting, $\tau_{i}=\sigma(\alpha\cdot(\frac{i}{20}-0.5)+\beta)$ with $\alpha=3$ and $\beta=0.5$, which we normalise into the range $\tau_{0}=0$ and $\tau_{1}=1$. This scheduler shifts integration steps at intermediate scales towards smaller scales, increasing the importance therein.
B.2.4 Censored flow matching for physical bounds
@@
Static thresholding, where the predictions are clipped into its physical bounds, can enforce these bounds [43]. However, models trained with the mean-squared error are biased and prone to spurious fluctuations near these boundaries [77]. These fluctuations can be particularly problematic in autoregressive predictions, where they may amplify instabilities in long-term simulations. Generative flow models that incorporate bounds via reflection or log-barriers [78, 79] enforce a bound-avoiding behaviour such that the bound is never reached, leading to unphysical predictions, where, e.g., all grid points exhibit sea ice, similar to a truncated distribution approximation. Instead, we introduce censored flow matching, accounting for static thresholding during training.
In censoring, all predictions that exceed a bound are exactly mapped to that bound. We generally assume that there exist true sea-ice dynamics without physical bounds, e.g., in the tropics, sea ice would still melt even though no sea ice is observed there. When the observed data is censored to a bound, its dynamics are unknown, and we have just the information that the dynamics exceeded the bound. In this case, the approximated velocity should increase the probability that the reproduced prediction exceeds the bound as well. To achieve this, we use the predicted velocity to estimate an approximated cumulative distribution function (CDF) from the variational Gaussian distribution. The CDF is then evaluated at the bounded velocity, and its logarithm replaces the mean-squared error loss for censored data points. A visualisation of the targetted training process is shown in Fig. 10.
Figure 10: Schematic showing the principle of censored flow matching. For two data cases (a and b), we draw independent noise during training and estimate a linear interpolant (dotted lines), connect the data and the noise. For a (blue), the data sample, $z_{1,a}$, is not on the bound (uncensored) and the targetted velocity points toward the data. For b (red), the data sample, $z_{1,b}$, is on the bound and, hence, censored; the true data value is unknown, it is only known that it lies between -$\infty$ and the bound as indicated by the red horizontal line. Here, the targetted velocity points towards an increase in the probability that the recovered data exceeds the bound.
In this framework, the variational Gaussian distribution, Equation (32), defines the latent distribution before the terminal state has been censored,
$q_{\boldsymbol{\theta}}(\widetilde{\mathbf{z}}_{1}\mid\mathbf{z}_{\tau})=\mathcal{N}\left(\widehat{\boldsymbol{\mu}}_{\boldsymbol{\theta}},\frac{1}{2}\Big{(}(1-\tau)\cdot\mathbf{s}(\tau)\Big{)}^{2}\mathbf{I}\right).$ (34)
@@
$q_{\boldsymbol{\theta}}(z_{1,k,l}\mid\mathbf{z}_{\tau})=\begin{cases}\frac{1}{s_{k}}\varphi(\frac{z_{1,k,l}-\mu_{\boldsymbol{\theta},k,l}}{s_{k}}),&z^{L}_{k,l}<\widetilde{z}_{1,k,l}<z^{U}_{k,l}\\ \Phi(\frac{z^{L}_{k,l}-\mu_{\boldsymbol{\theta},k,l}}{s_{k}}),&\text{if }\widetilde{z}_{1,k,l}=z^{L}_{k,l}\\ \Phi(\frac{\mu_{\boldsymbol{\theta},k,l}-z^{U}_{k,l}}{s_{k}}),&\text{if }\widetilde{z}_{1,k,l}=z^{U}_{k,l}\\ 0,&\text{otherwise},\end{cases}$ (36)
@@
$-\mathrm{I}(\widetilde{z}_{1,k,l}=z^{U}_{k,l})\left[\log\Phi\Big{(}\frac{\widehat{v}_{\boldsymbol{\theta},k,l}-u_{k,l}}{s_{k}}\Big{)}\right],$ (37)
@@
$\mathcal{L}_{\mathrm{GenSIM}}(\boldsymbol{\theta},\mathbf{s})=\mathbb{E}_{\tau,\mathbf{z}_{0},\mathcal{D}}\left[\sum^{n_{\mathrm{vars}}}_{k=1}\sum^{n_{\mathrm{grid}}}_{l=1}J(\mu_{\boldsymbol{\theta},k,l},s_{k},z_{1,k,l},z^{L}_{k,l},z^{U}_{k,l})\right],$ (38)
and $\mathrm{I}(\cdot)$ is the indicator function. The predicted velocity is no longer only used to estimate the dynamics but also to estimate the probability that the bound is exceeded [34]; the neural network is trained in a mixed classification-regression setting, also called Tobit model [80]. In settings with physical bounds, the resulting estimator for the velocities is consistent and gives an unbiased estimate of the true but only partially observed velocities [81]. As schematically shown in Fig. 10, the loss function pushes the velocity towards a prediction that exceeds the bound, when the observed data is exactly on the bound, while the velocity in the uncensored case remains the same.
@@
$\widehat{\mathbf{z}}_{1\mid\tau}=\mathrm{minimum}\Big{(}\mathrm{maximum}\big{(}\mathbf{z}_{\tau}+\tau\widehat{\mathbf{v}}_{\boldsymbol{\theta}},(\mathbf{\mathbf{x}^{L}}-\mathbf{x}_{t})/\boldsymbol{\sigma}_{\mathrm{clim}}\big{)},(\mathbf{\mathbf{x}^{U}}-\mathbf{x}_{t})/\boldsymbol{\sigma}_{\mathrm{clim}}\Big{)}$ (39)
with $\mathrm{minimum}(\cdot,\cdot)$ and $\mathrm{maximum}(\cdot,\cdot)$ as minimum and maximum function, respectively. By applying Equation (26), we obtain the velocity from the potentially thresholded projected state,
$\widehat{\mathbf{v}}_{\boldsymbol{\theta}}=\frac{\widehat{\mathbf{z}}_{1\mid\tau}-\mathbf{z}_{\tau}}{1-\tau}.$ (40)
@@
Appendix C Neural network architectures
We employ two different neural networks architectures: for GenSIM, we use a transformer that works on subdomains, while we use a U-ViT architecture [62, 34] for the deterministic model to efficiently process the full Arctic domain at once. Both architectures are described in the following.
C.1 Transformer architecture
The velocity for the generative flow is approximated by a neural network, which is trained to work on subdomains of size $80\times 80$. Our curvilinear mesh is a structured mesh with $512\times 512$ non-Cartesian cells, which are partially masked due to land. Accounting for these considerations, we built our neural network around a diffusion transformer architecture [40], inspired by design choices from large language models and from recent efficient diffusion models [82].
The neural network takes as main input $24$ channels–consisting of the six state variables $\mathbf{z}_{\tau}$, the six sea-ice state variables $\mathbf{x}_{t}$, and the twelve atmospheric forcing variables $\mathbf{F}_{t:t+12\,\textrm{h}}$ (four variables for two time steps and four channels for the degree day features). The transformer needs as additional input the mesh in Cartesian coordinates in a north polar stereographic projection, and the land mask ($1=\text{ocean}$, $0=\text{land}$). The scalar input that is converted into an embedding is comprised of the pseudo time $\tau\in[0,1]$, the resolution in kilometer, and three labels for the distribution augmentation [47, 35].
@@ C.1 Transformer architecture
The scalar inputs are converted into an embedding concurrently to the input tokenisation. The pseudo time, resolution, and labels are independently processed by combining random Fourier features (RFF)[83], which expand the number of features to $512$, with a linear layer, a Sigmoid Linear Unit (SiLU) activation function [84], and another linear layer, keeping the number of feature at $512$. At the end, each of the three branches is added together and activated with a SiLU to obtain the embedding.
The tokens are processed in a sequence of $8$ transformer blocks, in which self-attention blocks are interleaved with small multi-layer perceptrons (MLP). Each of the self-attention blocks and MLPs is added to a short skip connection that directly links the input to the output of each block [85]. All self-attention blocks have the same structure:
  • The input is normalised by root-mean-squared normalisation (RMSNorm) [86] where the dimension-wise scaling coefficient is a linear combination of the embedding [40].
@@
  • The queries and keys are rotated by a two-dimensional rotary positional embedding (RoPE) [87, 42] based on the tokenised mesh and scaled by the resolution. This efficiently implement an anisotropic relative distance dependency into the self-attention block.
  • After applying RoPE, multi-head self attention [41] is applied along the feature dimension, extracting $8$ different spatial features per block.
@@
  • A gated SiLU activation function [88] activates the inflated features, before the output is deflated once again to $512$ features by a linear projection without bias.
@@
To stabilise the transformer by imitating a U-Net-like architecture [89, 90], we implement long-skip connections, linking the output of the tokeniser to the input for the final head, linking the output of the first transformer block to the input of the last transformer block, and so forth. As linking function, we use a linear combination [82], which is dimension-wise modulated by the embedding.
In the end, also called the head, the tokens are normalised by RMSNorm and scaled with the embedding. To extract features that can contain sharp edges, we apply a rectified linear unit (ReLU) activation function after the normalisation. In the end, we linearly project the $512$ features per token to $24$ output features ($6\cdot 2\times 2$). Applying pixel shuffle [91], the output features are rearranged to form $6$ predicted velocity values per cell.
@@
The transformer has in total $28.1\times 10^{6}$ parameters, which is smaller than typical generative models for image generation [82] and for weather prediction [48]. In early experiments, we have seen that we could increase the number of layers and the feature dimensionality to around $500\times 10^{6}$ parameters without overfitting to improve the scores. Yet, for efficiency, we constrain the number of parameters. The model for the standard deviation adds $1.6\times 10^{6}$ parameters to the training of the generative flow.
C.2 U-ViT architecture
In addition to the transformer architecture for GenSIM, we use a U-ViT architecture for the deterministic model. This architecture combines U-Net like convolutional blocks [89] with a vision transformer [62] in its bottleneck. The convolutional blocks make it feasible to extract features on the pan-Arctic domain without subdomain decomposition. However, in contrast to self-attention, convolutional blocks are based on a fixed weighting, which assumes a local Cartesian mesh. We additionally have to modify the convolutions to handle the land mask [12, 34]. Throughout the U-ViT, prior to the application of convolutional layers and when skip connections are involved, the data is masked with the land mask, setting masked cells to zero: the convolutions are hence similar to partial convolutions without reweighting [92].
As for the transformer, the input data is masked, and the mask is added to the masked input data. A first convolutional layer (no bias, $1\times 1$ kernel size) extracts $64$ channels from the $18$ inputted predictors ($6\times$ sea-ice state, $4\times 2$ atmospheric forcings, and $4\times$ degree-day features). Concurrently, the neural network extract an embedding from the augmentation labels, as similarly done for the transformer.
Afterwards, $6$ downsampling blocks reduce the spatial dimensions from $512\times 512$ to $8\times 8$, while increasing the number of channels from $64$ to $512$. Each downsample block combines two ConvNeXt blocks [93] with a downsampling layer.
ConvNeXt blocks are inspired by transformer blocks: the initial convolutional layer (no bias, $7\times 7$ kernel size) extracts spatial information, which is non-linearly combined by the subsequent multi-layer perceptron (MLP). The MLP has nearly the same structure as for the transformer with its RMSNorm modulated by the embedding. The output from the MLP is added to the block input in a short-skip connection.
The downsampling layer is a RMSNorm modulated by the embedding and followed by a convolutional layer (no bias, $2\times 2$ kernel size, $2\times 2$ stride size). While reducing the spatial dimension four times, the convolutional layer doubles the number of channels starting from the $64\times 64$ resolution. The mask is concurrently coarse-grained by maximum pooling and a $2\times 2$ window, keeping cell values if at least a single cell is unmasked in the window.
After reducing the resolution to $8\times 8$ by the downsampling blocks, the bottleneck consists of $8$ transformer blocks as they are similarly used in the transformer architecture: the extracted features and the reduced mask are flattened in space, while the mesh is averaged to the $8\times 8$ resolution and flattened as well. The only difference between the transformer architecture and the bottleneck are the lower number of tokens in the bottleneck and its missing long skip connections.
The resolution of the features extracted by the bottleneck is increased by $6$ upsampling blocks, which include long skip connections from the downsampling blocks to retain high-resolution features. The upsampling blocks mirror the strategy from the downsampling blocks with an upsampling layer followed by two ConvNeXt blocks.
The upsampling layer normalises the incoming features by RMSNorm, modulated by the embedding, before the resolution is doubled in $x$- and $y$-direction. To double the resolution, we use a convolutional layer (no bias, $1\times 1$ kernel size) with pixel shuffle, as similarly used for the head in the transformer. Up to a resolution of $128\times 128$, the convolution halves the number of channels. The features with the increased resolution are combined with the features from the long-skip connection, outputted by the last ConvNeXt block from the downsampling block at the same resolution. As for the transformer, we linearly combine the features, where the channels-wise weighting stems from the embedding.
The head that uses the output from the last upsampling block works similar to the head of the transformer: a ReLU follows a normalisation with a modulated RMSNorm. As the head already works on the $512\times 512$ resolution, the last layer only combines the extracted features with a convolution ($1\times 1$ kernel size) into predictions.
The convolutional layers add parameters to the neural network, while the bottleneck has the same number of parameters as the self-attention and MLP blocks from the transformer. In total, the U-ViT architecture has $32.2\times 10^{6}$ parameters.

### v2

@@ Acknowledgments
An interesting case arises when we employ the model with a single step, mapping directly from $\tau=0$ to $\tau=1$: there, the latent state for $\tau=0$ contains no information about the prediction. Consequently, if a neural network is trained for that, it would learn to disregard the latent state as input and to target a mean prediction (?), similar to the deterministic model. In this case, the model effectively reduces to a deterministic prediction model.
The weighting in the loss function for the deterministic model and flow matching is determined by the climatological standard deviation. However, each of the six predicted variables is subject to distinct physical processes, leading to varying degrees of predictability, which might be not captured by their climatological weighting (?).
We account for this in a data-dependent weighting by altering the Gaussian assumption of variational flow matching, Equation S30, and making some covariances learnable. As a simplification and exploiting the mean-field approximation, we introduce a per-variable standard deviation that depends on the pseudo time, $\mathbf{s}(\tau)$, obtaining
$q_{\boldsymbol{\theta}}(\mathbf{z}_{1}\mid\mathbf{z}_{\tau})=\mathcal{N}\left(\widehat{\boldsymbol{\mu}}_{\boldsymbol{\theta}},\frac{1}{2}\Big((1-\tau)\cdot\mathbf{s}(\tau)\Big)^{2}\mathbf{I}\right),$ (S33)
@@
$\mathcal{L}_{\textrm{VFM}}(\boldsymbol{\theta},\mathbf{s})=\mathbb{E}_{\tau,\mathbf{z}_{0},\mathcal{D}}\left[\frac{1}{2\cdot\big(\mathbf{s}(\tau)\big)^{2}}\lVert\mathbf{u}-\widehat{\mathbf{v}}_{\boldsymbol{\theta}}(\mathbf{z},\tau,\mathbf{x}_{t},\mathbf{F}_{t:t+12h})\rVert^{2}_{2}+\log\big(\mathbf{s}(\tau)\big)\right],$ (S34)
which has a similar formulation for diffusion models (?). The squared standard deviation, $\big(\mathbf{s}(\tau)\big)^{2}$, converges to half of the expected per-variable mean-squared error of the velocity and adapts automatically to the predictability of different variables.
Note that, while the weighting in the loss function is learned, the output of the neural network is still scaled by the climatological standard deviation. This disentanglement between the scaling and the learned weighting becomes especially interesting when the neural network is trained across different datasets or resolutions: we can set a constant scaling of the neural network and learn different weightings per dataset (Fig. S10a). The weighting concentrates around the relevant scales, acting as gradient normalisation and stabilising the training (?).
This learnable weighting can reduce the need to tune the scheduler during training and prediction. We can also use the learned standard deviation to guide our inference scheduler (Fig. S10b). The larger the learned standard deviation, the larger the expected error for that scale, and importance sampling suggests that we should oversample these scales (?). Hence, we want to concentrate our pseudo time scheduler during prediction to the scales where the expected error is large, which are in our case the small scales. While these small scales might be unimportant for perceptual generation like images, they can become especially important when the model is auto-regressively cycled, as errors at small scales can amplify over time and impact larger scales. As prediction scheduler, we use a shifted sigmoid weighting, $\tau_{i}=\sigma(\alpha\cdot(\frac{i}{20}-0.5)+\beta)$ with $\alpha=3$ and $\beta=0.5$, which we normalise into the range $\tau_{0}=0$ and $\tau_{1}=1$. This scheduler shifts integration steps at intermediate scales towards smaller scales, increasing the importance therein.
@@
Static thresholding, where the predictions are clipped into its physical bounds, can enforce these bounds (?). However, models trained with the mean-squared error are biased and prone to spurious fluctuations near these boundaries (?). These fluctuations can be particularly problematic in autoregressive predictions, where they may amplify instabilities in long-term simulations. Generative flow models that incorporate bounds via reflection or log-barriers (?, ?) enforce a bound-avoiding behaviour such that the bound is never reached, leading to unphysical predictions, where, e.g., all grid points exhibit sea ice, similar to a truncated distribution approximation. Instead, we introduce censored flow matching, accounting for static thresholding during training.
In censoring, all predictions that exceed a bound are exactly mapped to that bound. We generally assume that there exist true sea-ice dynamics without physical bounds, e.g., in the tropics, sea ice would still melt even though no sea ice is observed there. When the observed data is censored to a bound, its dynamics are unknown, and we have just the information that the dynamics exceeded the bound. In this case, the approximated velocity should increase the probability that the reproduced prediction exceeds the bound as well. To achieve this, we use the predicted velocity to estimate an approximated cumulative distribution function (CDF) from the variational Gaussian distribution. The CDF is then evaluated at the bounded velocity, and its logarithm replaces the mean-squared error loss for censored data points. A visualisation of the targetted training process is shown in Fig. S11.
In this framework, the variational Gaussian distribution, Equation S33, defines the latent distribution before the terminal state has been censored,
$q_{\boldsymbol{\theta}}(\widetilde{\mathbf{z}}_{1}\mid\mathbf{z}_{\tau})=\mathcal{N}\left(\widehat{\boldsymbol{\mu}}_{\boldsymbol{\theta}},\frac{1}{2}\Big((1-\tau)\cdot\mathbf{s}(\tau)\Big)^{2}\mathbf{I}\right).$ (S35)
@@
$q_{\boldsymbol{\theta}}(z_{1,k,l}\mid\mathbf{z}_{\tau})=\begin{cases}\frac{1}{s_{k}}\varphi(\frac{z_{1,k,l}-\mu_{\boldsymbol{\theta},k,l}}{s_{k}}),&z^{L}_{k,l}<\widetilde{z}_{1,k,l}<z^{U}_{k,l}\\ \Phi(\frac{z^{L}_{k,l}-\mu_{\boldsymbol{\theta},k,l}}{s_{k}}),&\text{if }\widetilde{z}_{1,k,l}=z^{L}_{k,l}\\ \Phi(\frac{\mu_{\boldsymbol{\theta},k,l}-z^{U}_{k,l}}{s_{k}}),&\text{if }\widetilde{z}_{1,k,l}=z^{U}_{k,l}\\ 0,&\text{otherwise},\end{cases}$ (S37)
@@
$-\mathrm{I}(\widetilde{z}_{1,k,l}=z^{U}_{k,l})\left[\log\Phi\Big(\frac{\widehat{v}_{\boldsymbol{\theta},k,l}-u_{k,l}}{s_{k}}\Big)\right],$ (S38)
@@
$\mathcal{L}_{\mathrm{GenSIM}}(\boldsymbol{\theta},\mathbf{s})=\mathbb{E}_{\tau,\mathbf{z}_{0},\mathcal{D}}\left[\sum^{n_{\mathrm{vars}}}_{k=1}\sum^{n_{\mathrm{grid}}}_{l=1}J(\mu_{\boldsymbol{\theta},k,l},s_{k},z_{1,k,l},z^{L}_{k,l},z^{U}_{k,l})\right],$ (S39)
and $\mathrm{I}(\cdot)$ is the indicator function. The predicted velocity is no longer only used to estimate the dynamics but also to estimate the probability that the bound is exceeded (?); the neural network is trained in a mixed classification-regression setting, also called Tobit model (?). In settings with physical bounds, the resulting estimator for the velocity is consistent and gives an unbiased estimate of the true but only partially observed velocity (?). As schematically shown in Fig. S11, the loss function pushes the velocity towards a prediction that exceeds the bound, when the observed data is exactly on the bound, while the velocity in the uncensored case remains the same.
@@
$\widehat{\mathbf{z}}_{1\mid\tau}=\mathrm{minimum}\Big(\mathrm{maximum}\big(\mathbf{z}_{\tau}+\tau\widehat{\mathbf{v}}_{\boldsymbol{\theta}},(\mathbf{\mathbf{x}^{L}}-\mathbf{x}_{t})/\boldsymbol{\sigma}_{\mathrm{clim}}\big),(\mathbf{\mathbf{x}^{U}}-\mathbf{x}_{t})/\boldsymbol{\sigma}_{\mathrm{clim}}\Big)$ (S40)
with $\mathrm{minimum}(\cdot,\cdot)$ and $\mathrm{maximum}(\cdot,\cdot)$ as minimum and maximum function, respectively. By applying Equation S27, we obtain the velocity from the potentially thresholded projected state,
$\widehat{\mathbf{v}}_{\boldsymbol{\theta}}=\frac{\widehat{\mathbf{z}}_{1\mid\tau}-\mathbf{z}_{\tau}}{1-\tau}.$ (S41)
@@
Transformer architecture
The velocity for the generative flow is approximated by a neural network, which is trained to work on subdomains of size $80\times 80$. Our curvilinear mesh is a structured mesh with $512\times 512$ non-Cartesian cells, which are partially masked due to land. Accounting for these considerations, we built our neural network around a diffusion transformer architecture (?), inspired by design choices from large language models and from recent efficient diffusion models (?).
The neural network takes as main input $24$ channels–consisting of the six state variables $\mathbf{z}_{\tau}$, the six sea-ice state variables $\mathbf{x}_{t}$, and the twelve atmospheric forcing variables $\mathbf{F}_{t:t+12\,\textrm{h}}$ (four variables for two time steps and four channels for the degree day features). The transformer needs as additional input the mesh in Cartesian coordinates in a north polar stereographic projection, and the land mask ($1=\text{ocean}$, $0=\text{land}$). The scalar input that is converted into an embedding is comprised of the pseudo time $\tau\in[0,1]$, the resolution in kilometer, and three labels for the distribution augmentation (?, ?).
@@
The scalar inputs are converted into an embedding concurrently to the input tokenisation. The pseudo time, resolution, and labels are independently processed by combining random Fourier features (RFF) (?), which expand the number of features to $512$, with a linear layer, a Sigmoid Linear Unit (SiLU) activation function (?), and another linear layer, keeping the number of feature at $512$. At the end, each of the three branches is added together and activated with a SiLU to obtain the embedding.
The tokens are processed in a sequence of $8$ transformer blocks, in which self-attention blocks are interleaved with small multi-layer perceptrons (MLP). Each of the self-attention blocks and MLPs is added to a short skip connection that directly links the input to the output of each block (?). All self-attention blocks have the same structure:
  • The input is normalised by root-mean-squared normalisation (RMSNorm) (?) where the dimension-wise scaling coefficient is a linear combination of the embedding (?).
@@
  • The queries and keys are rotated by a two-dimensional rotary positional embedding (RoPE) (?, ?) based on the tokenised mesh and scaled by the resolution. This efficiently implement an anisotropic relative distance dependency into the self-attention block.
  • After applying RoPE, multi-head self attention (?) is applied along the feature dimension, extracting $8$ different spatial features per block.
@@
  • A gated SiLU activation function (?) activates the inflated features, before the output is deflated once again to $512$ features by a linear projection without bias.
@@
To stabilise the transformer by imitating a U-Net-like architecture (?, ?), we implement long-skip connections, linking the output of the tokeniser to the input for the final head, linking the output of the first transformer block to the input of the last transformer block, and so forth. As linking function, we use a linear combination (?), which is dimension-wise modulated by the embedding.
In the end, also called the head, the tokens are normalised by RMSNorm and scaled with the embedding. To extract features that can contain sharp edges, we apply a rectified linear unit (ReLU) activation function after the normalisation. In the end, we linearly project the $512$ features per token to $24$ output features ($6\cdot 2\times 2$). Applying pixel shuffle (?), the output features are rearranged to form $6$ predicted velocity values per cell.
@@
The transformer has in total $28.1\times 10^{6}$ parameters, which is smaller than typical generative models for image generation (?) and for weather prediction (?). In early experiments, we have seen that we could increase the number of layers and the feature dimensionality to around $500\times 10^{6}$ parameters without overfitting to improve the scores. Yet, for efficiency, we constrain the number of parameters. The model for the standard deviation adds $1.6\times 10^{6}$ parameters to the training of the generative flow.
Figure S1: The different regions used throughout the manuscript with the following names: 1=central Arctic; 2=Beaufort sea; 3=Chukchi sea; 4=east Siberian sea; 5=Laptev sea; 6=Kara sea; 7=Barents sea; 8=east Greenland sea; 9=Baffin bay, Labrador seas, and gulf of St. Lawrence; 10=Hudson bay; 11=Canadian archipelago. The black hatched rectangle is the region in the central Arctic used to estimate the spectra and fractal scaling. The regions are extracted from NSIDC’s regions on the EASE-Grid 2.0 with a 3.125 km resolution (?).
Figure S2: Overview of GenSIM’s short-term forecast qualities. Shown are: (a) the improvement in the root-mean-squared error (RMSE) of GenSIM’s ensemble mean compared to the deterministic model and the ratio between the ensemble standard deviation to the RMSE of GenSIM’s ensemble mean, both are averaged over all ice-covered grid points (sit=thickness, sic=concentration, sid=damage, siu=$u$-drift, siv=$v$-drift, snt=snow-on-ice thickness); (b) the 10-day RMSE improvement of GenSIM’s ensemble mean to the deterministic model split up into regions, averaged across all six predictive variables after normalizing with their climatology; (c) the 10-day spectrum of the sea-ice thickness in the central Arctic, averaged in log-space across all forecasts; (d) accuracy of the 10-day forecast for the sea-ice extent with the median accuracy for the ensemble members. The used regions for (b) and (c) are depicted in Fig. S1.
Figure S3: GenSIM’s ensemble mean has a lower RMSE than a deterministic baseline across regions, variables, and seasons.The score card, depicting the relative root-mean-squared error (RMSE) of GenSIM’s ensemble mean to the RMSE of the deterministic baseline for different regions, variables, and seasons, as estimated over all ice-covered grid points (sit=thickness, sic=concentration, sid=damage, siu=$u$-drift, siv=$v$-drift, snt=snow-on-ice thickness). A blue RMSE corresponds to an improvement of GenSIM, while a red one shows a degradation of the forecast. The regions (1–5) are depicted and explained in Fig. S1 and remaining regions are shown in Fig. S4.
Figure S4: Continuation of Fig. S3 for regions (6–11).
Figure S5: Fractal scaling of GenSIM’s dynamics. Triangulated total deformation rates as inferred from Sentinel-1 synthetic aperture radar (SAR) imagery (a), from advecting with neXtSIM’s predicted drift (b), and from advecting with GenSIM’s predicted drift (c) for the period 2018-02-16 03:00–2018-02-25 15:00. The deformation rates are estimated at the original data resolution and image pairs with an averaged signal-to-noise ratio <0.5<0.5 are filtered out. Note that the rates are temporally ordered with the newest images on top. (d) Distribution of total deformation rates for GenSIM and neXtSIM at the nominal resolution of around 12 km and an averaging window of 12 hours. (e) First three distributional moments of the total deformation rate (q) with coarse-grained spatial scale and the fitted power laws as linear lines for an averaging window of 12 hours. (f) First three distributional moments of the total deformation rate (q) with increasing averaging window at the nominal resolution of around 12 km. In (d)–(f), the total deformation rates are estimated based on continuously tracked particles from 2017-12-01 03:00 to 2018-03-31 15:$0$, which were advected with the drift of GenSIM and neXtSIM, respectively. The particles were initialised in a $65\times 65$ mesh at the cell centres of an area in the central Arctic. The power laws in (e) and (f) are fitted based on the relations $\left\langle\dot{\epsilon}^{q}_{\mathrm{tot}}(L)\right\rangle\sim L^{-\beta(q)}$ and $\left\langle\dot{\epsilon}^{q}_{\mathrm{tot}}(T)\right\rangle\sim T^{-\alpha(q)}$, respectively (?).
Figure S6: GenSIM predicts the slowing in the decrease of the sea-ice extent. (a, b) Monthly averaged sea-ice extent for April and September (a) and seasonal cycle of the sea-ice extent averaged for 2015–2018 (b), both inferred for satellite products from OSISAF and NSIDC, the PIOMAS reanalysis, and the simulations from neXtSIM and GenSIM. (c–e) Sea-ice extent trend over 20 years in the OSISAF product (c) and in the GenSIM simulation (d) for different months, as well as the September trend for all products.
Figure S7: GenSIM captures the evolution of sea-ice thickness and area. (a–b) Monthly averaged sea-ice thickness for April and September (a) and seasonal cycle averaged for 2015–2018 (b). (c–d) Monthly averaged sea-ice area for April and September (c) and seasonal cycle averaged for 2015–2018 (d). Please note that the CS2/SMOS and IceSAT2 product for the sea-ice thickness are only available for the winter months, while the OSISAF and NSIDC products for sea-ice area are available over the full years.
Figure S8: GenSIM simulates the decrease of the sea-ice age. (a–c) Sea-ice age for GenSIM (a) for neXtSIM (b), and as observed (c) at 2003-12-31 15:00 UTC. (d–f) Sea-ice age for GenSIM (d), for neXtSIM (e), and as observed (f) at 2018-12-31 15:00 UTC. The black contour lines indicate multi-year ice. Note that the contour lines for the observed age in (c) and (f) are smoothed with a $4\times 4$ averaging filter for visualisation purposes. The sea-ice age is post-processed from the simulated sea-ice drift and sea-ice concentration, while the observations are the EASE-Grid Sea Ice Age product from the NSIDC (?).
Figure S9: GenSIM performs similarly across resolutions. (a) Monthly averaged sea-ice volume on an EASE-100 km mesh for the PIOMAS reanalysis and simulations from GenSIM at the original resolution ($12\,\text{km}$) and at coarse resolution ($96\,\text{km}$). Note that the volume estimates are smaller than in Fig. 3 due to a coarse-grained masking. (b–d) Divergence rate for 2024-02-13 03:$0$ estimated with the drift from GenSIM at the original resolution (b), the drift coarse-grained from $12\,\text{km}$ to $96\,\text{km}$ (c), the drift from GenSIM at coarse resolution (d).
Figure S10: Learned weighting of the flow matching loss (a) The learned weighting, averaged across all six variables, at different resolutions indicates the importance of difference scales at different resolutions. (b) Used in an importance-sampling-like mechanism, the learned standard deviation can guide the tuning of the pseudo time scheduler during prediction. Here, the focus is on smaller scales, as opposed to a uniform scheduler.
Figure S11: Schematic showing the principle of censored flow matching. For two data cases (a and b), we draw independent noise during training and estimate a linear interpolant (dotted lines), connect the data and the noise. For a (blue), the data sample, $z_{1,a}$, is not on the bound (uncensored) and the targetted velocity points toward the data. For b (red), the data sample, $z_{1,b}$, is on the bound and, hence, censored; the true data value is unknown, it is only known that it lies between -$\infty$ and the bound as indicated by the red horizontal line. Here, the targetted velocity points towards an increase in the probability that the recovered data exceeds the bound.